Use the following information from separate companies a through f: Net Income (Loss) Interest Expense Income Taxes a. $ 142,000 $ 59,640 $ 35,500 b. 136,600 40,980 49,176 c. 145,100 10,157 69,648 d. 120,700 43,452 50,694 e. 92,300 15,691 35,074 f. (39,760 ) 83,894 0 Compute times interest earned.
Answer #1

Times interest earned for company A = (Net Income + Interest Expense + Income Taxes)/ Interest Expense

= (142,000 + 59,640 + 35,500)/59,640

= 237,140/59,640

= 3.98

Times interest earned for company B = (Net Income + Interest Expense + Income Taxes)/ Interest Expense

= (136,600 + 40,980 + 49,176)/40,980

= 226,756/40,980

= 5.53

Times interest earned for company C = (Net Income + Interest Expense + Income Taxes)/ Interest Expense

= (145,100 + 10,157 + 69,648)/10,157

= 224,905/10,157

= 22.14

Times interest earned for company D = (Net Income + Interest Expense + Income Taxes)/ Interest Expense

= (120,700 + 43,452 + 50,694)/43,452

= 214,846/43,452

= 4.94

Times interest earned for company E = (Net Income + Interest Expense + Income Taxes)/ Interest Expense

= (92,300 + 15,691 + 35,074)/15,691

= 143,065/15,691

= 9.12

Times interest earned for company F = (Net Income + Interest Expense + Income Taxes)/ Interest Expense

= (- 39,760 + 83,894 + 0)/83,894

= 44,134/83,894

= 0.53

Note: All the ratios have been rounded to two decimal places.

Exact answer may slightly differ due to rounding off.

Please ask if you have any query related to the question. Thank you